INFRARED PHOTODISSOCIATION SPECTROSCOPY OF SMALL, MASS-SELECTED PROTONATED WATER CLUSTERS

\underline{R. S. WALTERS} and M. A. DUNCAN, University of Georgia, Department of Chemistry, Athens, GA, 30602-2556.

H^{+}(H_{2}O)_{n} and H^{+}( H_{2}O)_{n}Ar_{m} complexes are produced in a laser vaporization pulsed nozzle source, mass-selected, and studied by infrared photodissociation spectroscopy. Photofragment yields are measured as a function of energy in the region near the symmetric and asymmetric stretching vibrations of water (3657 cm^{−}^{1}, 3756 cm^{−}^{1}). Dissociation occurs by the loss of intact water molecules in the pure complexes. The argon-tagged complexes fragment by the loss of Ar, producing larger fragment yields and sharper spectra than the corresponding pure clusters. The spectra of the smaller complexes (n=1-8) are compared to theory and previous experiments. At larger cluster sizes, the spectra simplify to a closely spaced doublet at the center of the O-H region (3700 cm^{−}^{1}) and a broad feature in the hydrogen bonded region. It is concluded that at a specific cluster size (n=10), most of the water molecules are included in the hydrogen bonded network, giving rise to only one type of O-H stretch.